Textra Plastics produces parts for a variety of small machine manufacturers. Most products go through two operations, molding and trimming, before they are ready for packaging. Expected costs and activities for the molding department and for the trimming department for 2017 follow.

Molding Trimming
Direct labor hours 42,000 DLH 52,000 DLH
Machine hours 38,500 MH 3,800 MH
Overhead costs $ 740,000 $ 570,000

  
Data for two special order parts to be manufactured by the company in 2017 follow.

Part A27C Part X82B
Number of units 11,000 units 51,500 units
Machine hours
Molding 5,800 MH 1,160 MH
Trimming 4,200 MH 1,050 MH
Direct labor hours
Molding 5,500 DLH 2,800 DLH
Trimming 700 DLH 3,500 DLH


Required
1. Compute the plantwide overhead rate using direct labor hours as the base.
2. Determine the overhead cost assigned to each product line using the plantwide rate computed in requirement 1.
Answer #1

1) Plantwide overhead rate = Estimated overhead/Estimated direct labor hours

= (740000+570000)/(42000+52000)

Plantwide overhead rate = $13.94 per labor hour

2) Overhead cost assigned to each product line

Part A27C = (5500+700)*13.94 = $86428

Part X82B = (2800+3500)*13.94 = $87822